A toy car rolls of the top of a table at a horizontal speed of 2.4m/s and falls to the floor in a time of 0.50 seconds. How tall is the table, and how far from the base of the table will the car land?
How tall answer 1.23 m
how far 1.20 m
could u show me the approach.. thanks!!!

Answer by Boreal(15235)

s=(1/2)at^2; a=9.8 m/sec^2, t=0.5 sec
s=(1/2)2.45=1.225 m
2.4 m/sec horizontally and hits after 0.5 sec.
The product is 1.2 m